iYuko (iyuko):

What is the value of x? Round your answer to the nearest tenth. x2 = 34 ±

iYuko (iyuko):

Do I square root 34?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

divide 34 by 2

iYuko (iyuko):

That wont work

OpenStudy (anonymous):

why not? its the right way to do it

iYuko (iyuko):

17*17 is no where close to 34

iYuko (iyuko):

Do 5.8*5.8

OpenStudy (anonymous):

but, isnt it x multiplied by 2?

iYuko (iyuko):

@Mateaus

OpenStudy (anonymous):

\[x^2 = 34\]\[x = \sqrt{34}\]

OpenStudy (anonymous):

Think of 2 numbers that you can use to multiply and get 34

iYuko (iyuko):

5.9

iYuko (iyuko):

But that is more thatn 34

OpenStudy (anonymous):

I mean. What two numbers multiplied give you 34? Example: 3*5 = 15

OpenStudy (anonymous):

if there is none, then you can't simplify further.

iYuko (iyuko):

WHat do you mean?

OpenStudy (anonymous):

You would leave it as \[x = \pm \sqrt{34}\]

OpenStudy (anonymous):

because there is no further simplification like the one from before where you could break it up into two.

iYuko (iyuko):

Thank you.